hello,I need help in kinds of numbers,converting ratio to a decimal,and converting each number to a fraction in lowest terms.
Hope you can help me:)

Answer by josmiceli(19441)

You can convert any decimal to a fraction
right away:
.1 = 1/10
.12 = 12/100
.123 = 123/1000
.1234 = 1234/10000
etc.
Suppose you have 7/56 for a fraction
If you do the same thing to both the top and the bottom
of a fraction, the value of it doesn't change. What if I
divide both top and bottom by 7 ?
(7/7) / (56/7)
7/7 is just 1, so now I have
1 / (56/7)
7 goes into 56 8 times, so the fraction is now 1/8
-----------
Now what if I multiplied the top and bottom of 1/8 by 125?
( 1*125) / (8*125)
125 / 1000
In decimal, this is .125, so
1/8 = .125
------------
hope this helps